OpenStudy (anonymous):

how do solve 4xover3/7over10x

OpenStudy (turingtest):

please rewrite w/ parentheses

OpenStudy (lalaly):

\[\huge{\frac{\frac{4x}{3}}{\frac{7}{10x}}}\]

OpenStudy (lalaly):

?

OpenStudy (turingtest):

remember: (a/b)/(c/d)=(a/b)(d/c) so in this case: (4x/3)/(7/10x)=(4x/3)(10x/7)=40x^2/21